OpenStudy (anonymous):

what is the result of adding 2.5x10^3 and 3.5x10^2

OpenStudy (anonymous):

2.5·10¹⁰ x 3.5·10-² = (2.5·3.5)·10^(10-2) = 8.75·10⁸

OpenStudy (aaronq):

2.5x10^3 = 2500 and 3.5x10^2 = 350